# On Discounted Infinite-Time Mean Field Games

### v2

@@ 6.1 Regularity of the FBSDE solutions
  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on ×$\mathbb{R}\times\mathcal{P}_{2}$.
While the validity of these two conditions has not been established for general cases, we can prove them for linear cases. If we set $b(x,\mu,a)=b_{1}x+b_{2}\bar{\mu}+b_{3}a$ and $f(x,\mu,a)=b_{4}x\bar{\mu}+Ax^{2}+Ca^{2}$, and we further assume that all assumptions in Theorem 4.4 and (4.27) are satisfied, the FBSDEs (6.1) and (6.2) become the following linear equations:
@@
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(b_{1}X_{t}^{\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left[(b_{1}-r)Y_{t}^{\xi}+2AX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.4)
@@
$\begin{cases}\mathrm{d}X_{t}^{x,\xi}=\left(b_{1}X_{t}^{x,\xi}-\frac{b^{2}_{3}}{2C}Y_{t}^{x,\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{x,\xi}=-\left[(b_{1}-r)Y_{t}^{x,\xi}+2AX_{t}^{x,\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right]\mathrm{d}t+Z_{t}^{x,\xi}\mathrm{d}B_{t},\\ X_{0}^{x,\xi}=x.\end{cases}$ (6.5)
We have already proved that for any initial state $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and $x\in\mathbb{R}$, equations (6.4) and (6.5) have unique solutions in $L^{2}_{r}$.
@@
For two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution, their corresponding solutions to (6.4) have the same expectations. That is
$\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[X_{t}^{\xi_{2}}],\quad~{}\mbox{for a.e.$t\geq 0.$}$ (6.6)
Proof. Fix $\xi_{1}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ and the corresponding solution $X^{\xi_{1}}_{t}$ to equation (6.4). Taking $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, we consider the following two infinite-time FBSDEs:
$\begin{cases}\mathrm{d}X_{t}^{\xi}=\left(pX_{t}^{\xi}-qY_{t}^{\xi}+b_{2}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}^{\xi}=-\left(uY_{t}^{\xi}+vX_{t}^{\xi}+b_{4}\mathbb{E}[X_{t}^{\xi}]\right)\mathrm{d}t+Z_{t}^{\xi}\mathrm{d}B_{t},\\ X_{0}^{\xi}=\xi.\end{cases}$ (6.7)
and
$\begin{cases}\mathrm{d}\tilde{X}_{t}^{\xi}=\left(p\tilde{X}_{t}^{\xi}-q\tilde{Y}_{t}^{\xi}+b_{2}\phi_{t}\right)\mathrm{d}t+\mathrm{d}B_{t},\\ \mathrm{d}\tilde{Y}_{t}^{\xi}=-\left(u\tilde{Y}_{t}^{\xi}+v\tilde{X}_{t}^{\xi}+b_{4}\phi_{t}\right)\mathrm{d}t+\tilde{Z}_{t}^{\xi}\mathrm{d}B_{t},\\ \tilde{X}_{0}^{\xi}=\xi.\end{cases}$ (6.8)
Where $p=b_{1},q=\frac{b^{2}_{3}}{2C},u=b_{1}-r,v=2A$, and according to the argument in Theorem 4.4, each of the aforementioned equations admits a unique solution in $L^{2}_{r}$. We denote the solutions for Equation (6.7) with initial states ($\xi_{1},\xi_{2}$) by ($X_{t}^{\xi_{1}},X_{t}^{\xi_{2}}$), and that for Equation (6.8) under identical initial states by ($\tilde{X}_{t}^{\xi_{1}},\tilde{X}_{t}^{\xi_{2}}$).
For Equation (6.8), let $P_{t}^{\xi}=m\tilde{X}_{t}^{\xi}-\tilde{Y}_{t}^{\xi}$ where $m$ is the positive solution of $qm^{2}-(p+u)m-v=0$. Then
$\begin{split}\mathrm{d}P_{t}^{\xi}&=\left[(mp+v)\tilde{X}_{t}^{\xi}-(mq-u)\tilde{Y}_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t}\\ &=\left[(mq-u)P_{t}^{\xi}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+(m-\tilde{Z}_{t}^{\xi})\mathrm{d}B_{t}.\end{split}$ (6.9)
This demonstrates that $P^{\xi}$ satisfies an infinite-time backward stochastic differential equation
$\mathrm{d}P_{t}=\left[(mq-u)P_{t}+(b_{2}m-b_{4})\phi_{t}\right]\mathrm{d}t+Z_{t}\mathrm{d}B_{t}$ (6.10)
and exhibits no dependence on $\xi$. Applying ([14], Theorem 4) with the facts that $mq-u\geq-u=r-b_{1}\geq\frac{r}{2}$ and $\phi_{t}\in L^{2}_{r}$, the above equation admits a unique solution $(P,Z)\in L^{2}_{r}$. Then we know $\tilde{X}_{t}^{\xi}$ satisfy the following SDE,
$\mathrm{d}\tilde{X}_{t}^{\xi}=\left[(p-mq)\tilde{X}_{t}^{\xi}+qP_{t}+b_{2}\phi_{t}\right]\mathrm{d}t+\mathrm{d}B_{t},\quad\quad\tilde{X}_{0}^{\xi}=\xi.$ (6.11)
It is easy to see that $\tilde{X}_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{2}}$ have the same distribution, so we have $\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}].$
Since $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]$, $X_{t}^{\xi_{1}}$ and $\tilde{X}_{t}^{\xi_{1}}$ are the same. Then we have $\phi_{t}=\mathbb{E}[X_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{1}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]$, which means $\tilde{X}_{t}^{\xi_{2}}$ is also a solution to FBSDE (6.7) with initial $\xi_{2}$. By the uniqueness of the solution, we get that $\mathbb{E}[{X}_{t}^{\xi_{2}}]=\mathbb{E}[\tilde{X}_{t}^{\xi_{2}}]=\mathbb{E}[X_{t}^{\xi_{1}}]$. Now we finish the proof.
Remark 6.2
For any two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution and $x\in\mathbb{R}$, the solutions $Y^{x,\xi_{1}},Y^{x,\xi_{2}}$ to (6.5) are the same. We can say that $Y^{x,\xi}_{0}$ depends depends solely on the distribution of $\xi$.
Lemma 6.3
For FBSDEs (6.4) and (6.5), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. Then we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
@@
$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\mathbb{E}\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.14)

### v3

@@ 6 Viscosity solution to distribution-dependent elliptic PDE
  • $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ is jointly continuous on $\mathbb{R}\times\mathcal{P}_{2}$.
6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
We shall prove the weak uniqueness for the infinite-time FBSDE:
$\begin{cases}\mathrm{d}X_{t}=B(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+\mathrm{d}B_{t},\\ \mathrm{d}Y_{t}=-F(t,X_{t},Y_{t},\mathcal{L}_{X_{t}})dt+Z_{t}\mathrm{d}B_{t}.\end{cases}$ (6.4)
We know that under Assumption 2.1, the above FBSDE has unique solution $(X_{t},Y_{t},Z_{t})$ in $L_{r}^{2}$ with initial $X_{0}$. We then give the definitions of strong and weak uniqueness.
Definition 6.1 (Strong uniqueness)
We say that the strong uniqueness holds for FBSDE (6.4) if on any admissible set-up $(\Omega,{\cal F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B)$, for any two $\mathbb{F}$-progressively measurable $L^{2}_{r}$ three-tuples
$(X^{1}_{t},Y^{1}_{t},Z^{1}_{t})_{t\geq 0},\quad(X^{2}_{t},Y^{2}_{t},Z^{2}_{t})_{t\geq 0}$ (6.5)
satisfying the FBSDE (6.4) with the same initial condition $X_{0}$ (up to an exceptional event), it holds that
$\mathbb{E}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|X^{1}_{t}-X^{2}_{t}|^{2}+|Y^{1}_{t}-Y^{2}_{t}|^{2}+|Z^{1}_{t}-Z^{2}_{t}|^{2}\right)\mathrm{d}t\right]=0$ (6.6)
Definition 6.2 (Weak uniqueness)
For any two set-ups $(\Omega^{1},\mathcal{F}^{1},\mathbb{P}^{1},\mathbb{F}^{1})$ and $(\Omega^{2},\mathcal{F}^{2},\mathbb{P}^{2},\mathbb{F}^{2})$ with inputs $(X_{0}^{1},B^{1})$ and $(X_{0}^{2},B^{2})$, $X_{0}^{1},X_{0}^{2}$ have the same law on $\mathbb{R}$. We say the weak uniqueness holds for FBSDE (6.4) if the $L^{2}_{r}$ solutions $(X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on corresponding set-ups have the same distribution.
We use the same scheme as the one developed by Yamada and Watanabe to prove that path-wise uniqueness of solutions of FBSDE’s implies uniqueness in the sense of probability law.
Theorem 6.3
Assume that on $(\Omega,\mathcal{F},\mathbb{P},\mathbb{F})$ with input $(X_{0},B_{t})_{t\geq 0}$, the FBSDE has a unique strong solution $(X_{t},Y_{t},Z_{t})$. Then the law of $(X_{0},B_{t},X_{t},Y_{t},\int_{0}^{t}Z_{s}\mathrm{d}s)$ only depends on $\mathcal{L}(X_{0})$.
Proof. At first we denote by $C([0,\infty);\mathbb{R})$ the space of continuous $\mathbb{R}$-valued functions on $0,\infty)$ equipped with the metric of uniform convergence on compacts:
$d(\omega^{1},\omega^{2})=\sum_{n\geq 0}2^{-n}\sup_{t\in[0,n]}\max(|\omega_{t}^{1}-\omega_{t}^{2}|,1).$ (6.7)
And define
$\begin{split}&\Omega_{\text{input}}\triangleq\mathbb{R}\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{output}}\triangleq C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R})\times C([0,\infty);\mathbb{R}),\\ &\Omega_{\text{canon}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}.\end{split}$ (6.8)
Let us consider two filtered probability spaces $(\Omega^{i},\mathcal{F}^{i},\mathbb{P}^{i},\mathbb{F}^{i}),i=1,2$. Denote by $Q^{1}$ and $Q^{2}$ the distribution of $(X_{0}^{1},B_{t}^{1},X_{t}^{1},Y_{t}^{1},\int_{0}^{t}Z^{1}_{s}\mathrm{d}s)_{t\geq 0}$ and $(X_{0}^{2},B_{t}^{2},X_{t}^{2},Y_{t}^{2},\int_{0}^{t}Z^{2}_{s}\mathrm{d}s)_{t\geq 0}$ on $\Omega_{\text{canon}}=\Omega_{\text{input}}\times\Omega_{\text{output}}$, by $Q_{\text{input}}$ the common distribution of the processes $(X_{0}^{1},B_{t}^{1}),(X_{0}^{2},B_{t}^{2})$ on $\Omega_{\text{input}}$.
Let us now define for $i\in\{1,2\}$,
$q^{i}(\omega_{\text{input}};F):\Omega_{\text{input}}\times\mathcal{B}(\Omega_{\text{output}})\rightarrow[0,1]$ (6.9)
as the regular conditional probability for $\mathcal{B}(\Omega_{\text{output}})$ given $\omega_{\text{input}}\in\Omega_{\text{input}}$ (under $Q^{i}$). It satisfies:
  • $\forall\omega_{\text{input}}\in\Omega_{\text{input}}$, $q^{i}(\omega_{\text{input}};\cdot)$ is a probability measure on $(\Omega_{\text{output}},\mathcal{B}(\Omega_{\text{output}}))$.
  • $\forall F\in\mathcal{B}(\Omega_{\text{output}})$, the mapping $(\xi,w)\rightarrow q^{i}(\xi,w;F)$ is ${\cal B}(\mathbb{R})\otimes{\cal B}(C([0,\infty),\mathbb{R}))$-measurable.
  • $\forall F\in{\cal B}(\Omega_{\text{output}}),\forall G\in{\cal B}(\Omega_{\text{input}})$:
    $Q^{i}(G\times F)=\int_{G}q^{i}(\omega_{\text{input}};F)Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}).$ (6.10)
Next, we need a enlarged space $(\Omega_{\text{total}},\mathcal{G},Q)$ to support all processes. Define
$\Omega_{\text{total}}\triangleq\Omega_{\text{input}}\times\Omega_{\text{output}}\times\Omega_{\text{output}},$ (6.11)
and $\mathcal{G}$ is the completion of the $\sigma$-field ${\cal B}(\Omega_{\text{canon}})\otimes{\cal B}(\Omega_{\text{output}})$ by the collection $\cal N$} of all null sets under the probability measure
$Q(G\times F_{1}\times F_{2})=\int_{G}q^{1}(\omega_{\text{input}};F_{1})q^{2}(\omega_{\text{input}};F_{2})Q_{\text{input}}(\mathrm{d}\omega_{\text{input}}),$ (6.12)
where $F_{1},F_{2}\in{\cal B}(\Omega_{\text{output}}),G\in{\cal B}(\Omega_{\text{input}})$.
We observe that $Q(G\times F_{1}\times\Omega_{\text{output}})=Q^{1}(G\times F_{1}),Q(G\times\Omega_{\text{output}}\times F_{2})=Q^{2}(G\times F_{2}).$ Inparticular, we denote by $(\xi,w,x^{1},y^{1},\zeta^{1},x^{2},y^{2},\zeta^{2})$ the canonical process on $\Omega_{\text{total}}$, then $(\xi,w,x^{1},y^{1},\zeta^{1})$ has distribution $Q^{1}$ and $(\xi,w,x^{2},y^{2},\zeta^{2})$ has distribution $Q^{2}$.
We define $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ by
$z^{i}_{t}=\begin{cases}\lim_{n\rightarrow\infty}n\left(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+}\right)\quad~\mbox{if the limit exists,}\\ 0\quad~\mbox{otherwise.}\end{cases}$ (6.13)
[…]
$\zeta^{i}_{t}=\int_{0}^{t}z^{i}_{s}\mathrm{d}s,\quad t\geq 0.$ (6.14)
Moreover, we have
$\begin{split}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|z^{i}_{t}|\mathrm{d}t&\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}\mathbb{E}^{Q}\int_{0}^{\infty}\mathrm{e}^{rt}|n(\zeta^{i}_{t}-\zeta^{i}_{(t-\frac{1}{n})+})|^{2}\mathrm{d}t\\ &=\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n^{2}\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}|\int_{(t-\frac{1}{n})+}^{t}Z_{s}^{i}\mathrm{d}s|^{2}\mathrm{d}t\\ &\leq\mathop{\underline{\rm lim}}_{n\rightarrow\infty}n\mathbb{E}^{\mathbb{P}^{i}}\int_{0}^{\infty}\mathrm{e}^{rt}\int_{(t-\frac{1}{n})+}^{t}|Z_{s}^{i}|^{2}\mathrm{d}s\mathrm{d}t\\ &\leq\mathbb{E}^{Q^{i}}\int_{0}^{\infty}\mathrm{e}^{rs}|Z_{s}^{i}|^{2}\mathrm{d}s,\end{split}$ (6.15)
the last inequality following from Fubini’s theorem.
Let us now endow $(\Omega_{\text{total}},{\cal G},Q)$ with the filtration $\mathbb{G}$, where $\mathbb{G}=\{{\cal G}_{t}\}_{t\geq 0}$ is the complete and right-continuous augumentation under $Q$ of the canonical filtration
${\cal H}_{t}=\{(\xi,w_{s},x^{1}_{s},y^{1}_{s},\zeta^{1}_{s},x^{2}_{s},y^{2}_{s},\zeta^{2}_{s});0\leq s\leq t\}$ (6.16)
on $\Omega_{\text{total}}$. It’s easy to see that $\xi$ is ${\cal G}_{0}$-measurable and that $(w_{s},x^{1}_{s},y^{1}_{s},z^{1}_{s},x^{2}_{s},y^{2}_{s},z^{2}_{s})_{t\geq 0}$ is $\{{\cal G}_{t}\}_{t\geq 0}$-progressively measurable. Moreover, for $i\in\{1,2\}$,
$Q\left\{\omega\in\Omega_{\text{total}};(x,w,x^{i},y^{i},\zeta^{i})\in A\right\}=Q^{i}\left\{(X_{0}^{i},B^{i},X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\right\};\quad A\in{\cal B}(\Omega_{\text{canon}}).$ (6.17)
Actually, we just have to prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion: we follow the proof given in ([6], Remark 1.6). Let us firstly define
$\pi_{t}:C([0,\infty);\mathbb{R})\rightarrow C([0,t];\mathbb{R}),\quad h\rightarrow h|_{[0,t]},$ (6.18)
and
$\pi^{\prime}_{t}:\gamma\rightarrow\gamma_{t},\quad h\rightarrow h|_{[0,t]},$ (6.19)
where $\gamma=\Omega_{\text{output}}$ and
$\gamma_{t}=C([0,t];\mathbb{R})\times C([0,t];\mathbb{R})\times C([0,t];\mathbb{R}).$ (6.20)
Endowing $C([0,t];\mathbb{R})$ and $\gamma_{t}$ with their borelian $\sigma$-fields, we define
${\cal K}_{t}\triangleq\sigma\{\pi_{t}\},\quad{\cal K}^{\prime}_{t}\triangleq\sigma\{\pi^{\prime}_{t}\}.$ (6.21)
Using the separability of the spaces $C([0,t];\mathbb{R})$ and $\gamma_{t}$, we see that
${\cal K}_{t}=\sigma\{w_{s};0\leq s\leq t\},$ (6.22)
and that $\forall i\in\{1,2\},\forall A\in{\cal K}^{\prime}_{t}$, the set $\{(X^{i},Y^{i},\int_{0}^{\cdot}Z^{i}_{s}\mathrm{d}s)\in A\}$ belongs to ${\cal F}^{i}_{t}$.
Now, considering $A\in{\cal K}^{\prime}_{t}$ , we want to show that, for $i\in\{1,2\}$, the map
$\Omega_{\text{input}}\rightarrow[0,1],\quad(\xi,w)\rightarrow q^{i}(\xi,w;A)$ (6.23)
is measurable with respect to the completion of the $\sigma$-field ${\cal B}(\mathbb{R})\otimes{\cal K}_{t}$ under the probability measure $Q_{\text{input}}$, denoted $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$. Indeed, let us consider $F\in{\cal B}(\mathbb{R})$, $G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then, $\forall i\in\{1,2\}$
@@ 6.1 Yamada-Watanabe theorem for infinite-time FBSDEs
$\begin{split}&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{G_{2}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\\ =&\mathbb{E}^{\mathbb{P}^{i}}\left[I_{F}(X_{0})I_{G_{1}}(B^{i})I_{A}(X^{i},Y^{i},\int_{0}^{\cdot}Z_{s}^{i}\mathrm{d}s)\right]\mathbb{E}^{\mathbb{P}^{i}}\left[I_{G_{2}}(B^{i})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{i}(\xi,w;A)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w).\end{split}$ (6.24)
Hence, the map $(\xi,w)\rightarrow q^{i}(\xi,w;A)$ is measurable with respect to $\overline{{\cal B}(\mathbb{R})\otimes{\cal K}_{t}}$.
Now we prove that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion. Let us consider $(A,A^{\prime})\in({\cal K}^{\prime}_{t})^{2},F\in{\cal B}(\mathbb{R}),G_{1}\in{\cal K}_{t}$ and $G_{2}\in\{w_{s}-w_{t};s\geq t\}$. Then,
@@
$\begin{split}&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\\ =&\int I_{F}(\xi)I_{G_{1}}(w)I_{G_{2}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\int I_{F}(\xi)I_{G_{1}}(w)q^{1}(\xi,w;A)q^{2}(\xi,w;A^{\prime})Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\int I_{G_{2}}(w)Q_{\text{input}}(\mathrm{d}\xi\mathrm{d}w)\\ =&\mathbb{E}^{Q}\left[I_{F}(\xi)I_{G_{1}}(w)I_{A}(x^{1},y^{1},\zeta^{1})I_{A^{\prime}}(x^{2},y^{2},\zeta^{2})\right]\mathbb{E}^{Q}\left[I_{G_{2}}(w)\right].\end{split}$ (6.25)
Noting that ${\cal H}^{t}={\cal B}(\mathbb{R})\otimes{\cal K}_{t}\otimes{\cal K}^{\prime}_{t}\otimes{\cal K}^{\prime}_{t}$, we conclude that $(w_{t})_{t\geq 0}$ is a $\{{\cal G}_{t}\}_{t\geq 0}$ Brownian motion.
At last, applying the same procedure on $(z^{i}_{t})_{t\geq 0},i\in\{1,2\}$ in ([5], Volume II, Lemma 1.27), we obtain that Qa.s.Q-a.s.: for all $0\leq t\leq T,i\in\{1,2\}$,
$\begin{cases}x^{i}_{t}=\xi+\int_{0}^{t}B(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds+w_{t},\\ y^{i}_{t}=y^{i}_{T}+\int_{t}^{T}F(s,x^{i}_{s},y^{i}_{s},\mathcal{L}_{x^{i}_{s}})ds-\int_{t}^{T}z^{i}_{s}\mathrm{d}w_{s},\\ \mathbb{E}^{Q}\left[\int_{0}^{\infty}\mathrm{e}^{-rt}\left(|x^{i}_{t}|^{2}+|y^{i}_{t}|^{2}+|z^{i}_{t}|^{2}\right)\mathrm{d}t\right]<\infty.\end{cases}$ (6.26)
Through the strong uniqueness, we know that under $Q$, the processes $(x^{1}_{t},y^{1}_{t},\zeta^{1}_{t})_{t\geq 0}$ and $(x^{2}_{t},y^{2}_{t},\zeta^{2}_{t})_{t\geq 0}$ have the same law, then we get the desired result.
Remark 6.4
For any two initial states $\xi_{1},\xi_{2}\in\mathbb{L}^{2}(\mathcal{F}_{0})$ with the same distribution and $x\in\mathbb{R}$, the solutions $Y^{x,\xi_{1}},Y^{x,\xi_{2}}$ to (6.2) are the same. We can say that $Y^{x,\xi}_{0}$ depends solely on the distribution of $\xi$.
6.2 Regularity of the FBSDE solutions
@@
For FBSDEs (6.1) and (6.2), we define $\mathcal{V}(x,\mu)\triangleq Y^{x,\xi}_{0}$ for some initial state $\xi$ with distribution $\mu$. Assuming all assumptions in Theorem 4.4 are satisfied, we have, for any $x,x^{\prime}\in\mathbb{R}$ and $\mu,\mu^{\prime}\in\mathcal{P}_{2}$, there exists a constant $C>0$, such that
@@
$\begin{split}|Y^{x_{1},\xi_{1}}_{0}-Y^{x_{2},\xi_{2}}_{0}|^{2}&\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+W_{2}^{2}(X^{\xi_{1}}_{t},X^{\xi_{2}}_{t})\right]\mathrm{d}t\\ &\leq C_{1}\mathbb{E}\int_{0}^{\infty}\mathrm{e}^{-rt}\left[(\hat{X}^{x,\xi}_{t})^{2}+(\hat{Y}^{x,\xi}_{t})^{2}+(\hat{X}^{\xi}_{t})^{2}\right]\mathrm{d}t.\end{split}$ (6.29)